# AVG-LLaVA: A Large Multimodal Model with Adaptive Visual Granularity

### v2

@@ Abstract
Recently, when dealing with high-resolution images, dominant large multimodal models (LMMs) usually divide them into multiple local images and one global image, which will lead to a large number of visual tokens. In this work, we introduce AVG-LLaVA, an LMM that can adaptively select the appropriate visual granularity based on the input image and instruction. This approach not only reduces the number of visual tokens and speeds up inference, but also improves the overall model performance. Specifically, we introduce the following modules based on LLaVA-NeXT: (a) a visual granularity scaler that includes multiple pooling layers to obtain visual tokens with different granularities; (b) a visual granularity router, which includes a Transformer layer, an MLP layer, and a voter layer, used to select the appropriate visual granularity based on the image and instruction. Furthermore, we propose RGLF, a novel training paradigm that aims at aligning the granularity predicted by the router with the preferences of the LMM, without the need for additional manually annotated data. Extensive experiments and analysis show that AVG-LLaVA achieves superior performance across 11 benchmarks, as well as significantly reduces the number of visual tokens and speeds up inference (e.g., an 85.3% reduction in visual tokens and a 2.53$\times$ increase in inference speed on the AI2D benchmark). The code will be available at https://github.com/DeepLearnXMU/AVG-LLaVA.
@@ 1 Introduction
Recently, the field of artificial intelligence (AI) has witnessed a significant advancement in large multimodal models (LMMs) (OpenAI, 2023b; Zhu et al., 2023; Dai et al., 2023; Liu et al., 2023a; 2024a), marking a further step toward artificial general intelligence (AGI). Most existing LMMs follow the structure of LLaVA (Liu et al., 2023a; 2024a), which includes a vision encoder to embed images into visual tokens and a connector to map them into the word embedding space. Subsequently, these visual tokens are fed into a large language model (LLM) (Touvron et al., 2023; OpenAI, 2023a; Chiang et al., 2023) for multimodal understanding and reasoning, alongside the word embeddings.
Figure 1: An example of VQA that requires information at different visual granularities. The image is from MSCOCO (Lin et al., 2014) dataset.
@@
Due to the limitations imposed by the fixed aspect ratio (e.g., 1:1) and low resolution (e.g., 336×336) used by visual encoders (e.g., CLIP-ViT (Radford et al., 2021)), earlier LMMs face challenges in processing high-resolution images with different aspect ratios. To deal with this limitation, dominant models, such as LLaVA-NeXT (Liu et al., 2024b), dynamically divide each input high-resolution image into multiple local images. These local images are encoded separately, and their tokens are then concatenated with the tokens of the original global image. This approach will lead to longer visual token sequences, such as 2880 visual tokens for a 672×672 image. However, in practice, such fine-grained visual information is not always necessary, and in some cases, coarse-grained visual information can even be more beneficial for model predictions. For instance, as shown in Figure 1, when the model is asked to recognize the number on the jersey, it requires relatively fine-grained visual information. In contrast, determining the color of the jersey only necessitates coarse-grained visual information.
@@
In this paper, we propose Adaptive Visual Granularity LLaVA (AVG-LLaVA), an LMM that can adaptively select the appropriate visual granularity based on the input image and instruction. The basic intuition behind our model is that humans only scrutinize images carefully when answering difficult questions; otherwise, a brief glance is sufficient. As displayed in Figure 2, AVG-LLaVA extends LLaVA-NeXT with a visual granularity scaler and a visual granularity router. The visual granularity scaler performs multiple rounds of pooling on visual tokens without training, each time halving the number of visual tokens, thus obtaining a series of visual features with different granularities. The visual granularity router adaptively selects the appropriate visual granularity features based on the input multi-granularity visual features and text features. By doing so, for images and instructions that do not require fine-grained details, the number of visual tokens can be reduced, which not only speeds up inference but also improves performance. This performance enhancement likely stems from the reduction of redundant information, as selecting the appropriate visual granularity makes it easier for the model to answer questions based on images effectively.
@@
Besides, we observe that it is challenging to train the visual granularity router directly through visual instruction tuning (Liu et al., 2023a). This may be because the router cannot learn the distinctions between different visual granularities from visual instruction tuning, making it difficult to learn how to select the most appropriate visual granularity based on the image and instruction. To deal with this issue, we propose a novel training paradigm called Ranking Granularity to align LMM Feedback (RGLF). This paradigm aligns router probabilities of multiple granularities with LMM preferences by a ranking loss (Hadsell et al., 2006; Hopkins & May, 2011; Liu et al., 2022), effectively aiding the router in distinguishing between different visual granularities and selecting the appropriate one.
We further evaluate AVG-LLaVA on 11 benchmarks including tasks from various types (e.g., general VQA and text-oriented VQA, etc.). Extensive experimental results show that AVG-LLaVA can effectively reduce the number of visual tokens and improve inference speed (e.g., an 85.3% reduction in visual tokens and a 2.53$\times$ increase in inference speed on the AI2D benchmark) while achieving better performance under the same base LLM.
@@ 2 Related Work
Large Multimodal Models
LLMs such as GPT-4 (OpenAI, 2023a), LLaMA (Touvron et al., 2023), and Gemini (Team et al., 2023) have achieved significant success in language understanding and generation. Benefiting from this, multimodal large models (LMMs) have garnered widespread attention. Flamingo (Alayrac et al., 2022), BLIP-2 (Li et al., 2023b), and LLaMA-adapter (Zhang et al., 2023) integrate a frozen visual encoder and trainable modules into a LLM, extending it into a LMMs. These models are then fine-tuned using plain image-text pairs, enabling them to process and perceive visual content. To further improve instruction-following abilities and response quality, LLaVA (Liu et al., 2023a) fine-tunes the entire model using visual instruction data generated by GPT-4. However, since these LMMs rely on CLIP-ViT to process images at a fixed resolution (e.g. 336x336), it hinders the LMMs from perceiving image details at higher resolutions.
High-Resolution LMMs
@@
Vision Token Reduction for LMMs
Recently, several methods are proposed to reduce the visual tokens for LMMs, including visual token pruning and merging. For example, CrossGET (Shi et al., 2023) introduces a cross-modal token for leveraging cross-modal information to make decisions on token selection and merging. LLaVA-PruMerge (Shang et al., 2024) employs the similarity between the class token and other tokens as a key criterion for pruning and merging vision tokens. Unlike the aforementioned methods, FastV (Chen et al., 2024b) finds that most image tokens receive inefficient attention after the second decoder layer, and thus prunes half of the image tokens. Similarly, VTW (Lin et al., 2024b) adopts a more aggressive strategy to prune all visual tokens at a certain layer. Unfortunately, despite the above methods effectively reducing the number of visual tokens, they often lead to a certain degree of decline in model performance. More recently, LLaVA-M3 (Cai et al., 2024) obtains multi-granularity visual features by merging visual tokens through pooling, enabling manual control of the tradeoff between inference cost and performance. Significantly different from these methods, our model can adaptively select the appropriate visual granularity based on the input image and instruction, improving model performance while reducing visual tokens.
@@ 3.1 Model Architecture
High-Resolution Image Encoding
@@
Given an input image $\mathbf{I}$in $\mathbb{R}^{H\times W\times 3}$, we follow common practice (Liu et al., 2024b) to divide it into multiple smaller local images $\mathbf{I}_{local}\in\mathbb{R}^{H_{v}\times W_{v}\times 3}$, where $H_{v}$ and $W_{v}$ are the resolution that the vision encoder is originally trained for. Then, these local images are individually encoded into a $H_{p}\times W_{p}$ grid of visual tokens $\mathbf{X}_{local}\in\mathbb{R}^{H_{p}\times W_{p}\times C}$ by the image encoder, where $C$ is the dimension of the visual encoder. To preserve the global context information of the input image, we resize the original image to ($H_{v}$, $W_{v}$) and encode it as global visual tokens. Finally, we map both global visual tokens and local visual tokens to the word embedding space through an MLP-based vision-language connector.
Visual Granularity Scaler
This module sequentially stacks 1$\times$2 and 2$\times$1 average pooling layers, thereby obtaining visual features at multiple granularities and preserving the spatial information. In this work, we consider CLIP-ViT-L-336 (Radford et al., 2021) as the visual encoder, and thus each image is encoded into 24$\times$24 grid of visual tokens. Then, these visual tokens are fed into the visual granularity scaler, obtaining visual tokens with a grid of 24$\times$12, 12$\times$12, 12$\times$6 and 6$\times$6, respectively. In this way, we can obtain visual tokens of different granularities in a fine-to-coarse manner without training.
Visual Granularity Router
@@
Different visual granularity features can be considered as different experts, so the Mixture of Experts (MoE) (Shazeer et al., 2017; Komatsuzaki et al., 2023; Lin et al., 2024a) structure is particularly well-suited for selecting the appropriate visual granularity. Unlike the previous MoE studies that use linear layers as routers, we propose a multi-layer structure as illustrated in Figure 2 to select the appropriate visual granularity based on the input image and the instruction. Specifically, when dealing with an image, we first flatten and concatenate its visual tokens of all granularities to form multi-granularity visual tokens $\mathbf{\overline{X}}_{v}=[\mathbf{X}_{v}^{1};\mathbf{X}_{v}^{2};...;\mathbf{X}_{v}^{N}]$, where $\mathbf{X}_{v}^{i}$ represents the visual tokens of the $i$-th granularity, and $N$ is the number of visual granularities [^1] To simplify the explanation, we use a single image as an example. In practice, we include a global image and multiple local images, and each image will go through the following steps. The final result will be obtained by averaging the results of all the images.. Then, these visual tokens are concatenated with the filtered instruction tokens $\mathbf{\overline{X}}_{instruct}$ to serve as the input for the visual granularity router. Here, $\mathbf{\overline{X}}_{instruct}$ is obtained by calculating the cosine similarity between the original instruction tokens $\mathbf{X}_{instruct}$ and the visual tokens with original granularity $\mathbf{X}_{v}$, retaining the top-k most relevant ones. Afterwards, we apply a single Transformer (Vaswani et al., 2017) layer to facilitate the fusion of visual tokens at different granularities with instruction tokens. Subsequently, an MLP is applied to each token to predict the appropriate visual granularity, resulting in the logits $\mathbf{Z}_{out}$in $\mathbb{R}^{L\times N}$, where $L$ is the number of both visual and instruction tokens. To vote for the most appropriate visual granularity, we use a learnable weight matrix (Voter) $\mathbf{W}$in $\mathbb{R}^{1\times L}$ to aggregate the logits predicted by all tokens, yielding the final logits $\mathbf{Z}_{final}\in\mathbb{R}^{1\times N}$. Finally, we use softmax to calculate the probability distribution of each visual granularity, where the visual tokens corresponding to the granularity with the highest probability are fed into the LLM.
@@ 3.2 Multi-stage Training
To effectively train our model, we carefully design a multi-stage training strategy, which consists of four stages, as illustrated in Figure 3.
Stage 1: Pretraining
During this stage, we only pretrain the vision-language connector on a plain image-caption dataset. Formally, we define the following cross-entropy loss for the next token prediction:
$\mathcal{L}_{1}=-\sum_{t=1}^{T}\mathrm{log}P(x_{t}|\mathbf{X}_{v},\mathbf{X}_{c,<t}),$ (1)
where $\mathbf{X}_{c,<t}$ are the caption tokens before the current prediction token $x_{t}$, and $T$ is the length of target text tokens. Note that we fix the parameters of the vision encoder and large language model, both of which have been pre-trained on large-scale data. Through this stage of training, the image features can be aligned with the pre-trained LLM word embeddings.
Stage 2: Visual Instruction Tuning
@@
In the second stage, we jointly train the visual encoder, vision-language connector, and LLM on high-quality visual instruction data. In this way, the LLM can be converted into an LMM, which is able to complete various multimodal tasks. Specifically, we perform next-token predictions with the following cross-entropy loss only on the answering part
$\mathcal{L}_{2}=-\sum_{t=1}^{T}\mathrm{log}P(x_{t}|\mathbf{X}_{v},\mathbf{X}_{instruct},\mathbf{X}_{a,<t}),$ (2)
@@
where $\mathbf{X}_{a}$ denotes the answer tokens.
Stage 3: Multi-Granularity Visual Instruction Tuning
Following the previous stages, we introduce the visual granularity scaler. As described in Section 3.1, this module does not contain trainable parameters and thus does not need to be trained. Therefore, we use the same data as in the stage 2 to train the visual encoder, vision-language connector, and LLM, enabling them to perceive and process visual features of $N$ different granularities. Formally, the loss at this stage is formulated as
$\mathcal{L}_{3}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{T}\mathrm{log}P(x_{t}|\mathbf{X}_{v}^{i},\mathbf{X}_{instruct},\mathbf{X}_{a,<t}).$ (3)
Figure 3: The overview of multi-stage training. Note that visual granularity scaler does not contain any parameters that require training.
Stage 4: Ranking Granularity to Align LMM Feedback
@@
Lastly, we introduce the visual granularity router into the model training, where all other modules are frozen, and only the router is trained. This stage allows the model to select the appropriate visual granularity based on the input image and instruction. Intuitively, a straightforward approach to training the router is to use the same visual instruction fine-tuning method as in previous stages. However, we find that the router trained with this method performs poorly. This could be due to the difficulty of visual instruction fine-tuning in effectively enabling the router to learn the differences between different visual granularities.
@@
To address the above issue, we propose RGLF, as illustrated in Figure 4, where the router is trained with a ranking loss, utilizing the feedback from the LMM fine-tuned with multi-granularity visual instructions as the ranking criterion. Concretely, for the given image and instructions, we let the LMM predict answers using visual tokens of different granularity $\mathbf{X}_{v}^{i}$ and calculate their respective log probabilities. Then, based on these log probabilities, we sort $\mathbf{X}_{v}^{1};\mathbf{X}_{v}^{2};...;\mathbf{X}_{v}^{N}$ in a descending order to obtain $\mathbf{X}_{\hat{v}}^{1};\mathbf{X}_{\hat{v}}^{2};...;\mathbf{X}_{\hat{v}}^{N}$. Given the visual tokens $\mathbf{X}_{\hat{v}}^{i}$ of the $i$-th granularity, we directly consider those tokens ($\mathbf{X}_{\hat{v}}^{1}$; $\mathbf{X}_{\hat{v}}^{2}$; …; $\mathbf{X}_{\hat{v}}^{i-1}$) ranked above it as positive examples and the remaining tokens ($\mathbf{X}_{\hat{v}}^{i+1}$; $\mathbf{X}_{\hat{v}}^{i+2}$; …; $\mathbf{X}_{\hat{v}}^{N}$) as negative ones. Afterwards, we use the router to give scores (log probability) $s_{i}$ for each $\mathbf{X}_{\hat{v}}^{i}$:
@@
where $\lambda_{ij}$ is the margin calculated as the difference in log probabilities between the answers predicted by the LLM using visual tokens of the $i$-th and $j$-th granularities:
$\lambda_{ij}=\frac{j-i}{|T|}\sum_{t=1}^{T}(\mathrm{log}P(x_{t}|\mathbf{X}_{\hat{v}}^{i},\mathbf{X}_{instruct},\mathbf{X}_{a,<t})-\mathrm{log}P(x_{t}|\mathbf{X}_{\hat{v}}^{j},\mathbf{X}_{instruct},\mathbf{X}_{a,<t})).$ (6)
When the preference of $\mathbf{X}_{v}^{j}$ is only slightly worse than $\mathbf{X}_{v}^{i}$, the margin will be small. Conversely, when $\mathbf{X}_{v}^{j}$ is significantly worse than $\mathbf{X}_{v}^{i}$, the margin will correspondingly increase. In this way, we can dynamically adjust the margin to obtain adaptively penalty degrees between different pairs.
@@
In summary, the total loss is defined as the weighted sum of two losses:
$\mathcal{L}_{4}=\mathcal{L}_{rank}+\alpha\mathcal{L}_{ce},$ (9)
where $\alpha$ is the hyperparameters used to maintain the balance between the ranking loss $\mathcal{L}_{rank}$ and cross-entropy loss $\mathcal{L}_{ce}$.
Figure 4: The overview of RGLF. Visual tokens of each granularity are concatenated with instruction tokens and then processed by the LMM to estimate the corresponding rewards. Visual granularity router optimizes the score (log probability) of each granularity based on the feedback from the LMM.

### v3

# AVG-LLaVA: An Efficient Large Multimodal Model with Adaptive Visual Granularity

@@ Abstract
[…]
AVG-LLaVA: An Efficient Large Multimodal Model with
Adaptive Visual Granularity
Zhibin Lan1,3thanks:   Work was done when Zhibin Lan was interning at Pattern Recognition Center, WeChat AI, Tencent Inc, China., Liqiang Niu2, Fandong Meng2, Wenbo Li1,3, Jie Zhou2, Jinsong Su1,3,4thanks:   Corresponding author. 1School of Informatics, Xiamen University, China, 2Pattern Recognition Center, WeChat AI, Tencent Inc, China, 3Key Laboratory of Digital Protection and Intelligent Processing of Intangible Cultural Heritage of Fujian and Taiwan (Xiamen University), Ministry of Culture and Tourism, China, 4Shanghai Artificial Intelligence Laboratory, China lanzhibin@stu.xmu.edu.cn, jssu@xmu.edu.cn
@@ 1 Introduction
Recently, the field of artificial intelligence (AI) has witnessed a significant advancement in large multimodal models (LMMs) (OpenAI, 2023b; Zhu et al., 2023; Dai et al., 2023; Liu et al., 2023b, 2024a), marking a further step toward artificial general intelligence (AGI). Most existing LMMs follow the structure of LLaVA (Liu et al., 2023b, 2024a), which includes a vision encoder to embed input images into visual tokens and a connector to map them into the word embedding space. Subsequently, these visual tokens are fed into a large language model (LLM) (Touvron et al., 2023; OpenAI, 2023a; Chiang et al., 2023) for multimodal understanding and reasoning (Li et al., 2023b; Liu et al., 2023a; Zhang et al., 2024; Lin et al., 2025), alongside the word embeddings.
Figure 1: An example of VQA from MSCOCO (Lin et al., 2014). Notably, responding to Question 1 necessitates fine-grained visual information, whereas responding to Question 2 requires only coarse-grained visual information.
@@
Due to the limitations imposed by the fixed aspect ratio (e.g., 1:1) and low resolution (e.g., 336×336) used by visual encoders (e.g., CLIP-ViT (Radford et al., 2021)), early LMMs face challenges in processing high-resolution images with different aspect ratios. To deal with this limitation, dominant models, such as LLaVA-NeXT (Liu et al., 2024b), dynamically divide each input high-resolution image into multiple local images. These local images are encoded separately, and their tokens are then concatenated with the tokens of the original global image. This approach will lead to longer visual token sequences, such as 2880 visual tokens for a 672×672 image. However, in practice, such fine-grained visual information is not always necessary, and in some cases, coarse-grained visual information can even be more beneficial for model predictions. For instance, as shown in Figure 1, when the model is asked to recognize the number on the jersey, it requires relatively fine-grained visual information. In contrast, determining the color of the jersey only necessitates coarse-grained visual information.
In this paper, we propose Adaptive Visual Granularity LLaVA (AVG-LLaVA), an LMM that can adaptively select the appropriate visual granularity based on the input image and instruction. The basic intuition behind our model is that humans only scrutinize images carefully when answering difficult questions; otherwise, a brief glance is sufficient.
@@
As displayed in Figure 2, AVG-LLaVA extends LLaVA-NeXT with a visual granularity scaler and a visual granularity router. The visual granularity scaler performs multiple rounds of pooling on visual tokens, each time halving the number of visual tokens, thus obtaining a series of visual features with different granularities. The visual granularity router adaptively selects the appropriate visual granularity features based on the input multi-granularity visual features and text features. By doing so, for images and instructions that do not require fine-grained details, the number of visual tokens can be reduced, which not only speeds up inference but also may improves performance. This performance enhancement likely stems from the reduction of redundant information, as selecting appropriate visual granularity makes it easier for the model to answer questions based on images effectively.
@@
Besides, we observe that it is challenging to train the visual granularity router directly through visual instruction tuning (Liu et al., 2023b). This may be because the router cannot learn the distinctions between different visual granularities from visual instruction tuning, making it difficult to learn how to select the most appropriate visual granularity based on the image and instruction. To deal with this issue, we propose a novel training paradigm called Ranking Granularity based on LMM Feedback (RGLF). This paradigm aligns router probabilities of multiple granularities with LMM preferences by a ranking loss (Hopkins and May, 2011; Liu et al., 2022), effectively aiding the router in distinguishing between different visual granularities and selecting the appropriate one.
[…]
@@ 2 Related Work
High-Resolution LMMs.
Large language models (LLMs) such as GPT-4 (OpenAI, 2023a), LLaMA (Touvron et al., 2023), and Gemini (Team et al., 2023) have achieved significant success in language understanding and generation, driving the development of LMMs that integrate vision encoders with LLMs and leverage visual instruction data for fine-tuning. However, early LMMs (Li et al., 2023b; Zhang et al., 2023; Liu et al., 2023b) rely on fixed-resolution (e.g., 336×336) CLIP-ViT to process images, which limits their ability to capture high-resolution image details.
@@
Vision Token Reduction for LMMs.
Recently, several methods are proposed to reduce the visual tokens for LMMs, including visual token compression and pruning. For example, LLaVA-UHD (Guo et al., 2024) adopts a QFormer-like (Dai et al., 2023) structure to compress visual tokens, while MG-LLaVA (Zhao et al., 2024) employs a convolution layer to compress high-resolution visual features. In contrast, CrossGET (Shi et al., 2023) introduces a cross-modal token for leveraging cross-modal information to make decisions on token selection and merging. LLaVA-PruMerge (Shang et al., 2024) employs the similarity between the class token and other tokens as a key criterion for pruning and merging vision tokens.
Furthermore, FastV (Chen et al., 2024b) finds that most image tokens receive inefficient attention after the second decoder layer, and thus prunes half of the image tokens. Similarly, VTW (Lin et al., 2024b) adopts a more aggressive strategy to prune all visual tokens at a certain layer. Unfortunately, despite the above methods effectively reducing the number of visual tokens, they often lead to a certain degree of decline in model performance. More recently, LLaVA-M3 (Cai et al., 2024) obtains multi-granularity visual features by merging visual tokens through pooling, enabling manual control of the tradeoff between inference cost and performance.
Significantly different from aforementioned methods, our model is a dynamic neural network (Han et al., 2022) that can adaptively select the appropriate visual granularity based on the input image and instruction, improving model performance while reducing the number of visual tokens.
@@ 3.1 Model Architecture
High-Resolution Image Encoding.
@@
Given an input image $\mathbf{I}$ $\in$ $\mathbb{R}^{H\times W\times 3}$, we follow common practice (Liu et al., 2024b) to divide it into multiple smaller local images $\mathbf{I}_{local}\in\mathbb{R}^{H_{v}\times W_{v}\times 3}$. Here, $H_{v}$ and $W_{v}$ are the resolution that the vision encoder is originally trained for. Then, these local images are individually encoded into a $H_{p}\times W_{p}$ grid of visual tokens $\mathbf{X}_{local}\in\mathbb{R}^{H_{p}\times W_{p}\times C}$ by the image encoder, where $C$ is the dimension of the visual encoder. To preserve the global context information of the input image, we resize the original image to $H_{v}\times W_{v}$) and encode it as global visual tokens. Finally, we map both global and local visual tokens to the word embedding space through an MLP-based vision-language connector.
Visual Granularity Scaler.
[…]
Visual Granularity Router.
Different visual granularity features can be considered as different experts, so the Mixture of Experts (MoE) (Shazeer et al., 2017; Komatsuzaki et al., 2023; Lin et al., 2024a; Zhang et al., 2025) structure is particularly well-suited for selecting the appropriate visual granularity. Unlike the previous MoE studies that use linear layers as routers, we propose a multi-layer structure as illustrated in Figure 2 to select the appropriate visual granularity based on the input image and the instruction.
@@
Specifically, when dealing with an image, we first flatten and concatenate its visual tokens of all granularities to form multi-granularity visual tokens $\mathbf{\overline{X}}_{v}=[\mathbf{X}_{v}^{1};\mathbf{X}_{v}^{2};...;\mathbf{X}_{v}^{N}]$, where $\mathbf{X}_{v}^{i}$ represents the visual tokens of the $i$-th granularity, and $N$ is the number of visual granularities [^1] To simplify the explanation, we use a single image as an example. In practice, we include a global image and multiple local images, and each image will go through the following steps. The final result will be obtained by averaging the results of all the images.. Then, these visual tokens are concatenated with the filtered instruction tokens $\mathbf{\overline{X}}_{instruct}$ to serve as the input for the visual granularity router. Here, $\mathbf{\overline{X}}_{instruct}$ is obtained by calculating the cosine similarity between the original instruction tokens $\mathbf{X}_{instruct}$ and the visual tokens with original granularity $\mathbf{X}_{v}$, retaining the top-k most relevant ones. Afterwards, we apply a single Transformer (Vaswani et al., 2017) layer to facilitate the fusion of visual tokens at different granularities with instruction tokens. Subsequently, an MLP is applied to each token for predicting the appropriate visual granularity, resulting in the logits $\mathbf{Z}_{out}$ $\in$ $\mathbb{R}^{L\times N}$, where $L$ is the number of both visual and instruction tokens. To vote for the most appropriate visual granularity, we use a learnable weight matrix (Voter) $\mathbf{W}$ $\in$ $\mathbb{R}^{1\times L}$ to aggregate the logits predicted by all tokens, yielding the final logits $\mathbf{Z}_{final}\in\mathbb{R}^{1\times N}$. Finally, we use softmax to calculate the probability distribution of visual granularities, where the visual tokens corresponding to the granularity with the highest probability are fed into the LLM.
@@ 3.2 Multi-stage Training
We provide a detailed description of the training procedures for AVG-LLaVA, which consists of two stages. The first stage endows the model with the ability to perceive and process multi-granularity visual information, while the second stage enables the model to select the appropriate granularity based on the image and instructions.
Stage 1: Multi-Granularity Visual Instruction Tuning.
@@
In this stage, we use the high-quality visual instruction data to train the visual encoder, vision-language connector, and LLM, enabling them to perceive and process visual features of $N$ different granularities. Specifically, we perform next-token prediction using visual features of different granularities and apply the cross-entropy loss only to the answering part, formulated as
$\mathcal{L}_{1}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{T}\mathrm{log}P(x_{t}|\mathbf{X}_{v}^{i},\mathbf{X}_{instruct},\mathbf{X}_{a,<t}),$ (1)
@@
where $\mathbf{X}_{a}$ are the answer tokens before the current prediction token $x_{t}$, and $T$ is the length of answer tokens.
Stage 2: Ranking Granularity Based on LMM Feedback.
@@
Then, we introduce the visual granularity router into the model training, where all other modules are frozen, and only the router is trained. This stage allows the model to select the appropriate visual granularity based on the input image and instruction. Intuitively, a straightforward approach to training the router is visual instruction fine-tuning. However, we find that the router trained with this method performs poorly. This could be due to the difficulty of visual instruction fine-tuning in effectively enabling the router to learn the differences between different visual granularities.
Figure 3: The overview of RGLF. Visual tokens of each granularity are concatenated with instruction tokens and then processed by the LMM to estimate the corresponding rewards. Visual granularity router optimizes the granularity selection based on the feedback from the LMM.
@@
To address the above issue, we propose RGLF, as illustrated in Figure 3, where the router is trained with a ranking loss, utilizing the feedback from the LMM fine-tuned with multi-granularity visual instructions as the ranking criterion. Concretely, for the given image and instructions, we let the LMM predict answers using visual tokens of different granularity and calculate their respective log probabilities. Then, based on these log probabilities, we sort $\mathbf{X}_{v}^{1};\mathbf{X}_{v}^{2};...;\mathbf{X}_{v}^{N}$ in a descending order to obtain $\mathbf{X}_{\hat{v}}^{1};\mathbf{X}_{\hat{v}}^{2};...;\mathbf{X}_{\hat{v}}^{N}$. Given the visual tokens $\mathbf{X}_{\hat{v}}^{i}$ of the $i$-th granularity, we directly consider those tokens ($\mathbf{X}_{\hat{v}}^{1}$; $\mathbf{X}_{\hat{v}}^{2}$; …; $\mathbf{X}_{\hat{v}}^{i-1}$) ranked above it as positive examples and the remaining tokens ($\mathbf{X}_{\hat{v}}^{i+1}$; $\mathbf{X}_{\hat{v}}^{i+2}$; …; $\mathbf{X}_{\hat{v}}^{N}$) as negative ones. Afterwards, we use the router to give scores (log probability) $s_{i}$ for each $\mathbf{X}_{\hat{v}}^{i}$:
@@
where $\lambda_{ij}$ is the log probability difference between the answers predicted by the LLM using visual tokens of the $i$-th and $j$-th granularities:
$\begin{split}\lambda_{ij}=\frac{j-i}{|T|}&\sum_{t=1}^{T}(\mathrm{log}P(x_{t}|\mathbf{X}_{\hat{v}}^{i},\mathbf{X}_{instruct},\mathbf{X}_{a,<t})\\ &-\mathrm{log}P(x_{t}|\mathbf{X}_{\hat{v}}^{j},\mathbf{X}_{instruct},\mathbf{X}_{a,<t})).\end{split}$ (4)
When the preference of $\mathbf{X}_{\hat{v}}^{j}$ is only slightly worse than $\mathbf{X}_{\hat{v}}^{i}$, the margin will be small. Conversely, when $\mathbf{X}_{\hat{v}}^{j}$ is significantly worse than $\mathbf{X}_{\hat{v}}^{i}$, the margin will correspondingly increase. In this way, we can dynamically adjust the margin to obtain adaptively penalty degrees between different pairs.
@@
Finally, the total loss is defined as the weighted sum of two losses:
$\mathcal{L}_{2}=\mathcal{L}_{rank}+\alpha\mathcal{L}_{ce},$ (7)
where $\alpha$ is the hyperparameter used to maintain the balance between the ranking loss $\mathcal{L}_{rank}$ and cross-entropy loss $\mathcal{L}_{ce}$.